Two table were sold at rs.990 each, one at a profit of 10%. What are the cost prices of each of the tables? In this sale, did the seller make a profit or loss? What is the percentage of profit or loss? ( Hint: In the first case, 110% of cost price is 90% of cost price is 990)
Answer by Maths68(1474)

Two table were sold at rs.990 each, one at a profit of 10%. What are the cost prices of each of the tables? In this sale, did the seller make a profit or loss? What is the percentage of profit or loss? ( Hint: In the first case, 110% of cost price is 90% of cost price is 990)
Let
original price of table sold at 10% profit = x
110% of x = 990
110/100 * x = 990
11/10 *x = 990
x = 990 * 10/11
x = 9900/11 = 900

Original Price of the table sold at 10% = 900
but
the selling price was 990 = 990-900 = Rs. 90 (Profit)

Total Price of two tables = 990 + 990 = 1980
Original price of other table = 1980-900 = 1080
But
selling price was 990 = 1080-990 = 90 (loss)
Profit - loss = 90-90 = 0
Therefore seller broke even.